I am very confused on using the substitution and elimination method on linear equations. Can you please explain it to me?
Answer by mananth(16946)

I will solve a simple problem by both methods. See if you can walk through
Elimination method
2 x + 1 y = 20 .............1

3 x -4 y = 8 .............2
Eliminate y
multiply (1)by 4
Multiply (2) by 1
8 x 4 y = 80
3 x -4 y = 8
Add the two equations
11 x = 88
/ 11
x = 8
plug value of x in (1)
2 x + 1 y = 20
16 + y = 20
y = 20 -16
y = 4
y = 4
Substitution Method

2 x + 1 y = 20 .............1
1 y = 20 + -2 x
/ 1 =
y = 20 -2 x
3 x -4 y = 8 .....................2
Substitute y in (2)
3 x -4 ( 20 + -2 x ) = 8
3 x -80 + 8 x = 8
3 x 8 x = 88
11 x = 88
/ 11
x= 8
Plug the value of x in (1)
2 x + 1 y = 20
2 * 8 + 1 y = 20
16 + 1 y = 20
1 y = 4
/ 1
y= 4
m.ananth@hotmail.ca